OpenStudy (anonymous):

Please, help me solve the proportion below. 3/x = x/4 x=

OpenStudy (luigi0210):

cross multiply

OpenStudy (anonymous):

3x = 4x

OpenStudy (anonymous):

actually... 12 = x^2, right?

OpenStudy (anonymous):

yes 3x=4x when you cross mutiply your awnser

OpenStudy (luigi0210):

no it's x^2=12

OpenStudy (anonymous):

okay... but that can't be the final answer

OpenStudy (luigi0210):

\[x^2=12\] \[x=\pm \sqrt{4*3}\] \[x=\pm2\sqrt{3}\]

OpenStudy (anonymous):

thanks, that was the correct answer :)

OpenStudy (luigi0210):

No problem ^_^